# Distilling Vision-Language Models on Millions of Videos

### v1

@@ Abstract
The recent advance in vision-language models is largely attributed to the abundance of image-text data. We aim to replicate this success for video-language models, but there simply is not enough human-curated video-text data available. We thus resort to fine-tuning a video-language model from a strong image-language baseline with synthesized instructional data. The resulting video-language model is then used to auto-label millions of videos to generate high-quality captions. We show the adapted video-language model performs well on a wide range of video-language benchmarks. For instance, it surpasses the best prior result on open-ended NExT-QA by 2.8%. Besides, our model generates detailed descriptions for previously unseen videos, which provide better textual supervision than existing methods. Experiments show that a video-language dual-encoder model contrastively trained on these auto-generated captions is 3.8% better than the strongest baseline that also leverages vision-language models. Our best model outperforms state-of-the-art methods on MSR-VTT zero-shot text-to-video retrieval by 6%.
@@ 2 Related Work
Video-language models can be adapted from image-laungage models given that image-based foundation models are pre-trained on web-scale image data. VideoCLIP [66] leverages a pre-trained CLIP model [47] as a frame-level feature extractor and fine-tunes video and text transformers on video datasets. VideoCoCa [68] builds on CoCa [72] and fine-tunes some temporal pooler layers to reason over time. Another line of research focuses on parameter efficient tuning, which is first shown effective on language modeling [29]. AIM [70] adapts pre-trained image models for efficient video understanding by freezing pre-trained weights and tuning a few lightweight adapters. Furthermore, to solve more complex video-language tasks like captioning and QA, researchers leverage the powerful LLMs as a universal interface and adapt LLMs to consume visual tokens. FrozenBiLM [69] leverages a frozen bi-directional language model for video QA. VideoChat [32] and VideoChatGPT [37] propose a chatbot-like interface to analyze video input. However, VideoChat only shows qualitative analysis while VideoChatGPT relies on a GPT-4 for quantitative evaluation, leading to inconsistency over time. LaViLa [79] develops a video-language model that densely narrates for a video. However, training the narrator assumes videos to be partially annotated. Our work takes a further step and shows that the adapted video-language model generalizes to million-scale unseen videos.
@@ 3 Preliminaries and Notations
Image-based VLMs take as input an image and a text sequence, which is often called a prompt [4] or an instruction [63], and outputs another text sequence that follows the prompt. Specifically, let ${\mathbf{x}}\in\mathbb{R}^{H\times W\times 3}$ denote an input image with height $H$ and width $W$, ${\mathbf{y}}=(s_{1},\cdots,s_{L_{i}})\in\{0,1\}^{{L_{i}}\times|\mathbb{S}|}$ the instruction, and ${\mathbf{z}}=(z_{1},\cdots,z_{L_{o}})\in\{0,1\}^{{L_{o}}\times|\mathbb{S}|}$ the output text that are tokenized [28] into sequences of discrete symbols. Here, $\mathbb{S}$ denotes the vocabulary set, and $L_{i}$ and $L_{o}$ are the sequence lengths of the instruction and output, respectively.
@@ 5.1 Datasets
Table 1: Dataset summary. ADP is short for adapting VLMs while CP is for contrastive pre-training a dual-encoder model. The evaluation tasks include text-to-video retrieval (TVR), action classification (CLS), video captioning (CAP), and video question answering (QA).
@@ 5.2 Harnessing the Distilled Pseudo-Captions
We harness and evaluate the distilled pseudo-captions for million-scale web-scraped videos, ${}^{+}$ and ${}^{+}$, using a dual-encoder model [47]. The model’s video understanding performance is a solid indicator of the pseudo-captions’ quality, and we show that they are of higher quality than the original text provided in VideoCC and InternVid.
@@ 5.3 Main Results
7777707070707007007007007,00070007{,}0007 , 0003535353540404040Number of used data in VideoCC (×$\times 10^{3}$)ptMSR-VTT R@1
@@
Distilled captions for video understanding. We continue to evaluate the distilled pseudo-captions by the corresponding dual-encoder model’s zero-shot performance on text-to-video retrieval and video classification. From Table 2, we see that the pseudo-captions distilled from our VLM significantly improve the dual-encoder over the original text in VideoCC and InternVid. On VideoCC, with all other settings being the same, the dual-encoder model trained on ${}^{+}$, achieves 48.2% Recall@1 on MSR-VTT, 11.2% better than the one trained on the original Alt-text. It also clearly surpasses the recent ViCLIP trained on InternVid, which contains 2×2$\times$2 × more unique videos than VideoCC. On InternVid, our model trained on ${}^{+}$ is 3.8% better than the baseline trained on the original InternVid’s auto-generated captions. It is worth noting that our adapted VLM is also “lighter-weight” compared to the multi-scale captioning pipeline in InternVid [61], which relies on both image captioning models (BLIP-2) [31] on multiple frames and an LLM to put them together. We also highlight the zero-shot top-1 and top-5 classification accuracy on Kinetics-600. For instance, the dual-encoder model trained on ${}^{+}$${}^{+}$ improves the baselines on VideoCC/InternVid by 12.5/2.0% top-1 accuracy.
Interestingly, we notice that the model trained on ${}^{+}$ works better on action recognition, while the one trained on ${}^{+}$ is better on video retrieval. This is probably because the InternVid videos are specifically collected based on action phrases [61], while VideoCC is seeded from image-captioning data [41]. Since the two datasets are complementary, combining them indeed leads to performance gains as shown in the last row in Table 2.
@@ 5.4 Ablation Studies
How should we do visual adaptation? We study several ways for visual adaptation in Table 5. The first option, i.e. freezing the visual encoder $F_{V}$ while fine-tuning the language model $F_{L}$, takes inspiration from LiT [77]. This leads to a drop of 1.1 CIDEr score compared to our default recipe, where $F_{V}$ is fine-tuned and $F_{L}$ frozen. We ascribe it to the visual discrepancy between images and videos: The downstream tasks in [77, 10] are mostly still images, the same as the large-scale pre-training data. In contrast, the videos of our interests have unique characteristics such as object movement, camera motion, and the resultant visual degradation. We also observe a performance drop if we fine-tune both $F_{V}$ and $F_{L}$. This recipe may be prone to over-fitting because the video-text dataset lacks diversity and quantity. Finally, we show that self-training with VideoCC pseudo-captions (details in Appendix E) improves captioning results by 1.2 CIDEr score, reaching 43.5. It is worth noting that this number is on par with the best-performing PaLI-X [8] which has $\times1$ more parameters and takes 2×2$\times$2 × more frames as input than ours.
@@ Appendix A Instruction-Following Templates
Temporal reasoning You are an AI visual assistant that can analyze a single video. You receive a few sentences, each describing the same video you are observing. The task is to use the provided caption, create a plausible question about the video, and provide the answer in detail.
Always answer as if you are directly looking at the video.
[…]
  Causal reasoning You are an AI visual assistant that can analyze a single video. You receive a few sentences, each describing the same video you are observing. The task is to use the provided caption, create a plausible question about the video, and provide the answer in detail.
Always answer as if you are directly looking at the video.
[…]
  Short QAs You are an AI visual assistant that can analyze a single video. You receive a few sentences, each describing the same video you are observing. The task is to use the provided caption, create a plausible question about the video, and provide a short answer with less than three words.
Always answer as if you are directly looking at the video.
[…]
Question: What did the baby girl on the left side wearing a grey t-shirt do with the egg after she is carrying it? Answer: The girl throws the egg at the head of the man.   Captions: A boy wearing black shorts is standing on the side of the swimming pool over small rocks and then he performs a backflip and injured himself. In the background, there is a swimming pool, rocks, trees, and people’s voices and water splashing sound is audible. Question: What was the boy wearing black shorts doing before he performing a backflip? Answer: The boy is standing on the side of the swimming pool over small rocks. Then he performs a backflip and injured himself.   Captions: A boy wearing a black t-shirt rides a black bicycle in a backward direction and falls on a gray surface. A girl wearing a black cloth is moving on a gray surface, stops, and looks back toward the boy. In the background, there are gray surfaces, buildings, metallic barrier poles, trees, boats, and the sky is visible, and sounds of people speaking and the wind are audible. Question: What did the girl do after the boy wearing a black t-shirt rides a black bicycle in a backward direction and falls on a gray surface? Answer: The girl wearing a black cloth is moving on a gray surface. After the boy falls on a gray surface, she stops and looks back toward the boy.
Captions: A baby girl on the left side wearing a grey t-shirt is carrying an egg then she throws the egg at the head of the man, then the egg falls on the ground and it breaks on a grey surface. A man wearing a red t-shirt sitting on his knees is talking with the baby girl on a grey surface. In the background, there is a grey car, a grey surface, a brown mat, and people speaking and crying sounds are audible. Question: Why did the egg fall on the ground and breaks? Answer: The egg was thrown by the girl at the head of the man sitting on his knees. However, it did not hit the man. Therefore, the egg falls on the ground and breaks on a grey surface.   Captions: A boy wearing black shorts is standing on the side of the swimming pool over small rocks and then he performs a backflip and injured himself. In the background, there is a swimming pool, rocks, trees, and people’s voices and water splashing sound is audible. Question: How did the boy standing on the side of the swimming pool over small rocks injure himself? Answer: The boy standing on the side of the swimming pool over small rocks perform a backflip. Instead of falling into the water, he hit on the small rocks. That is why he injured himself.   Captions: A boy wearing a black t-shirt rides a black bicycle in a backward direction and falls on a gray surface. A girl wearing a black cloth is moving on a gray surface, stops, and looks back toward the boy. In the background, there are gray surfaces, buildings, metallic barrier poles, trees, boats, and the sky is visible, and sounds of people speaking and the wind are audible. Question: How did the boy wearing a black t-shirt fall on a gray surface? Answer: The boy wearing a black t-shirt rides a black bicycle in a backward direction. Riding a black bicycle backward is slow and hard to balance. Also, the boy is not good at riding a black bicycle backward. As a result, he fell on a gray surface.
Captions: A baby girl on the left side wearing a grey t-shirt is carrying an egg then she throws the egg at the head of the man, then the egg falls on the ground and it breaks on a grey surface. A man wearing a red t-shirt sitting on his knees is talking with the baby girl on a grey surface. In the background, there is a grey car, a grey surface, a brown mat, and people speaking and crying sounds are audible. Question: who throws the egg at the man Answer: baby girl   Captions: A boy wearing black shorts is standing on the side of the swimming pool over small rocks and then he performs a backflip and injured himself. In the background, there is a swimming pool, rocks, trees, and people’s voices and water splashing sound is audible. Question: what kind of pool is in the background Answer: swimming   Captions: A boy wearing a black t-shirt rides a black bicycle in a backward direction and falls on a gray surface. A girl wearing a black cloth is moving on a gray surface, stops, and looks back toward the boy. In the background, there are gray surfaces, buildings, metallic barrier poles, trees, boats, and the sky is visible, and sounds of people speaking and the wind are audible. Question: what happens when the man loses control Answer: falls down
@@ D.2 Training the Dual-Encoder Model
Appendix E Self-training with Pseudo-captioned Videos
@@
Appendix F Comparing with human-labeled data
@@
In this section, we compare the performance of the dual-encoder model trained on the human-labeled data and pseudo-captions. We train a dual-encoder model on the human-labeled S-MiT because (1) it is the largest human-labeled video-caption dataset to date and (2) our video-language model that is used to generate pseudo-captions for unlabeled videos is trained on S-MiT first. The zero-shot retrieval and recognition performance is reported in Table 11 in comparison with the result on ${}^{+}$ and ${}^{+}$. We can see that the dual-encoder model trained on both ${}^{+}$ and ${}^{+}$ clearly outperforms the one trained on S-MiT in terms of MSR-VTT zero-shot text-to-video retrieval recall. This indicates that our adapted video-language model not only distills human-labeled video dataset, but also generalizes to unseen videos that are within the same domain. When looking at the retrieval result on VATEX and classification result on Kinetics-600, the dual-encoder model trained on either ${}^{+}$ or ${}^{+}$, however, is slightly inferior to that on S-MiT. We ascribe this to the semantic correlation between S-MiT and Kinetics/VATEX: S-MiT is built on top of Moments-in-Times (MiT) whose videos are all tagged with one action or activity label similar to the way Kinetics and VATEX is constructed and the action concepts between MiT and Kinetics are closely related.

### v2

@@ Abstract
The recent advance in vision-language models is largely attributed to the abundance of image-text data. We aim to replicate this success for video-language models, but there simply is not enough human-curated video-text data available. We thus resort to fine-tuning a video-language model from a strong image-language baseline with synthesized instructional data. The resulting video model by video-instruction-tuning (VIIT) is then used to auto-label millions of videos to generate high-quality captions. We show the adapted video-language model performs well on a wide range of video-language benchmarks. For instance, it surpasses the best prior result on open-ended NExT-QA by 2.8%. Besides, our model generates detailed descriptions for previously unseen videos, which provide better textual supervision than existing methods. Experiments show that a video-language dual-encoder model contrastively trained on these auto-generated captions is 3.8% better than the strongest baseline that also leverages vision-language models. Our best model outperforms state-of-the-art methods on MSR-VTT zero-shot text-to-video retrieval by 6%. As a side product, we generate the largest video caption dataset to date.
@@ 2 Related Work
Video-language models can be adapted from image-language models given that image-based foundation models are pre-trained on web-scale image data. VideoCLIP [67] leverages a pre-trained CLIP model [48] as a frame-level feature extractor and fine-tunes video and text transformers on video datasets. VideoCoCa [69] builds on CoCa [73] and fine-tunes some temporal pooler layers to reason over time. Another line of research focuses on parameter efficient tuning, which is first shown effective on language modeling [30]. AIM [71] adapts pre-trained image models for efficient video understanding by freezing pre-trained weights and tuning a few lightweight adapters. Furthermore, to solve more complex video-language tasks like captioning and QA, researchers leverage the powerful LLMs as a universal interface and adapt LLMs to consume visual tokens. FrozenBiLM [70] leverages a frozen bi-directional language model for video QA. VideoChat [33] and VideoChatGPT [38] propose a chatbot-like interface to analyze video input. However, VideoChat only shows qualitative analysis while VideoChatGPT relies on a GPT-4 for quantitative evaluation, leading to inconsistency over time. LaViLa [80] develops a video-language model that densely narrates for a video. However, training the narrator assumes videos to be partially annotated. Our work takes a further step and shows that the adapted video-language model generalizes to million-scale unseen videos.
@@ 3 Preliminaries and Notations
Image-based VLMs take as input an image and a text sequence, which is often called a prompt [4] or an instruction [64], and outputs another text sequence that follows the prompt. Specifically, let ${\mathbf{x}}\in\mathbb{R}^{H\times W\times 3}$ denote an input image with height $H$ and width $W$, ${\mathbf{y}}=(s_{1},\cdots,s_{L_{i}})\in\{0,1\}^{{L_{i}}\times|\mathbb{S}|}$ the instruction, and ${\mathbf{z}}=(z_{1},\cdots,z_{L_{o}})\in\{0,1\}^{{L_{o}}\times|\mathbb{S}|}$ the output text that are tokenized [29] into sequences of discrete symbols. Here $\mathbb{S}$ denotes the vocabulary set, and $L_{i}$ and $L_{o}$ are the sequence lengths of the instruction and output, respectively.
@@ 5.1 Datasets
Table 1: Dataset summary. ADP is short for adapting VLMs while CP is for contrastive pre-training a dual-encoder model. Evaluation tasks include text-to-video retrieval (TVR), action classification (CLS), video captioning (CAP), and video question answering (QA).
@@ 5.2 Harnessing the Distilled Pseudo-Captions
We harness and evaluate the distilled pseudo-captions for million-scale web-scraped videos, VideoCC+ and InternVid+, using a dual-encoder model [48]. The model’s video understanding performance is a solid indicator of the pseudo-captions’ quality, and we show that they are of higher quality than the original text in VideoCC and InternVid.
@@ 5.3 Main Results
7777707070707007007007007,00070007{,}0007 , 0003535353540404040Number of used data in VideoCC (×$\times 10^{3}$)MSR-VTT R@1Alt-text [42]Img. cap. + LLMVid. cap. (Ours)
@@
Distilled captions for video understanding. We continue to evaluate the distilled pseudo-captions by the corresponding dual-encoder model’s zero-shot performance on text-to-video retrieval and video classification. From Table 2, we see that the pseudo-captions distilled from our VLM significantly improve the dual-encoder over the original text in VideoCC and InternVid. On VideoCC, with all other settings being the same, the dual-encoder model trained on VideoCC+, achieves 48.2% Recall@1 on MSR-VTT, 11.2% better than the one trained on the original Alt-text. It also clearly surpasses the recent ViCLIP trained on InternVid, which contains 2×2$\times$2 × more unique videos than VideoCC. On InternVid, our model trained on InternVid+ is 3.8% better than the baseline trained on the original InternVid’s auto-generated captions. It is worth noting that our adapted VLM is also “lighter-weight” compared to the multi-scale captioning pipeline in InternVid [63], which relies on both image captioning models (BLIP-2) [32] on multiple frames and an LLM to put them together. We also highlight the zero-shot top-1 and top-5 classification accuracy on Kinetics-600. For instance, the dual-encoder model trained on VideoCC+/InternVid+ improves the baselines on VideoCC/InternVid by 12.5/2.0% top-1 accuracy.
Interestingly, we notice that the model trained on InternVid+ works better on action recognition, while the one trained on VideoCC+ is better on video retrieval. This is probably because the InternVid videos are specifically collected based on action phrases [63], while VideoCC is seeded from image-captioning data [42]. Since the two datasets are complementary, combining them indeed leads to performance gains as shown in the last row in Table 2.
@@ 5.4 Ablation Studies
How should we do visual adaptation? We study several ways for visual adaptation in Table 5. The first option, i.e. freezing the visual encoder $F_{V}$ while fine-tuning the language model $F_{L}$, takes inspiration from LiT [78]. This leads to a drop of 1.1 CIDEr score compared to our default recipe, where $F_{V}$ is fine-tuned and $F_{L}$ frozen. We ascribe it to the visual discrepancy between images and videos: The downstream tasks in [78, 9] are mostly still images, the same as the large-scale pre-training data. In contrast, the videos of our interests have unique characteristics such as object movement, camera motion, and the resultant visual degradation. We also observe a performance drop if we fine-tune both $F_{V}$ and $F_{L}$. This recipe may be prone to over-fitting because the video-text dataset lacks diversity and quantity. Finally, we show that self-training with VideoCC pseudo-captions (details in §E.3) improves captioning results by 1.2 CIDEr score, reaching 43.5. It is worth noting that this number is on par with the best-performing PaLI-X [10] which has $\times1$ more parameters and takes 2×2$\times$2 × more frames as input than ours.
@@
More ablations. We leave more ablations and discussions in the supplementary materials.
@@ Appendix B Examples of Video Captioning
Additionally, we show some representative statistics from the natural language understanding perspective in Table 7. We analyze the sentence length and frequency of unique verbs/nouns using Spacy [25]. Our generated video captions (VideoCC+) are nearly 50% longer than alt-text. They also contain more unique verbs and nouns than the original alt-text, indicating higher diversity.
@@ C.1 Data for Evaluating the Dual-Encoder Model
Average length # of uniq. verbs # of uniq. nouns
 VideoCC 10.66 8,000 13,097
VideoCC+ 15.74 8,317 29,712
Table 7: Statistics of original alt-text and pseudo-captions.
@@ D.2 Training the Dual-Encoder Model
Appendix E More ablations
E.1 Justifying the Two-Stage Adaptation Design
We choose the adaptation order based on the information flow of the system, i.e. the Vision-Language Model needs to first perceive well in order to speak correctly. Thus we first adapt visual encoder to videos and then instruct-tune language model for details and completeness. We conduct experiments of reversing the order in Table 9 and see a noticeable drop of WUPS (29.5$\rightarrow$17.9).
We also compare prompt-tuning (PT) [30], an adapter-based fine-tuning method during language adaptation. Though the captioning result is close, the zero-shot QA accuracy is much worse. This indicates that adapter methods are not aimed at general-purpose multi-modal models.
S-MiT (CIDEr) NExT-QA (WUPS)
 V $\rightarrow$ L 42.3 29.5
L $\rightarrow$ V 41.9 17.9
V $\rightarrow$ L (PT) 40.9 4.88
Table 9: Effect of different adaptation orders.
E.2 Effect of Data Scale for Adaptation
We first describe the criteria of choose training datasets in our adaptation. The goal of visual adaptation is to align video features with language, especially actions that hardly exist in images. This calls for unambiguous and clean video-text pairs. The goal of language adaptation is to learn a general-purpose video-language model from videos with detailed text annotations, i.e. narratives. We thus choose the best available datasets to serve these two goals: S-MiT is the largest human-annotated video captioning dataset; VidLN is the largest and most diverse video dataset with detailed human narratives.
We study the effect of the dataset size in Table 10 and Table 11. Following the motivations above, we measure the visual adaptation performance by supervised video captioning and the language adaptation performance by zero-shot QA. For both stages, our model benefits from more data. This justifies the scaling ability of our model.
% of S-MiT videos 0 1% 10% 100%
 S-MiT (CIDEr) 11.5 37.6 39.6 42.3
Table 10: Effect of data size in visual adaptation.
% of VidLN videos 0 1% 10% 100%
 NExT-QA (WUPS) 1.2 8.8 16.0 29.5
Table 11: Effect of data size in language adaptation.
E.3 Self-training with Pseudo-captioned Videos
@@
E.4 Comparing with human-labeled data
@@
In this section, we compare the performance of the dual-encoder model trained on the human-labeled data and pseudo-captions. We train a dual-encoder model on the human-labeled S-MiT because (1) it is the largest human-labeled video-caption dataset to date and (2) our video-language model that is used to generate pseudo-captions for unlabeled videos is trained on S-MiT first. The zero-shot retrieval and recognition performance is reported in Table 8 in comparison with the result on VideoCC+ and InternVid+. We can see that the dual-encoder model trained on both VideoCC+ and InternVid+ clearly outperforms the one trained on S-MiT in terms of MSR-VTT zero-shot text-to-video retrieval recall. This indicates that our adapted video-language model not only distills human-labeled video dataset, but also generalizes to unseen videos that are within the same domain. When looking at the retrieval result on VATEX and classification result on Kinetics-600, the dual-encoder model trained on either VideoCC+ or InternVid+, however, is slightly inferior to that on S-MiT. We ascribe this to the semantic correlation between S-MiT and Kinetics/VATEX: S-MiT is built on top of Moments-in-Times (MiT) whose videos are all tagged with one action or activity label similar to the way Kinetics and VATEX is constructed and the action concepts between MiT and Kinetics are closely related.
{mdframed}
[align=center,linewidth=1pt] Temporal reasoning
You are an AI visual assistant that can analyze a single video. You receive a few sentences, each describing the same video you are observing. The task is to use the provided caption, create a plausible question about the video, and provide the answer in detail.
Always answer as if you are directly looking at the video.
[…]
Causal reasoning
You are an AI visual assistant that can analyze a single video. You receive a few sentences, each describing the same video you are observing. The task is to use the provided caption, create a plausible question about the video, and provide the answer in detail.
Always answer as if you are directly looking at the video.
[…]
Short QAs
You are an AI visual assistant that can analyze a single video. You receive a few sentences, each describing the same video you are observing. The task is to use the provided caption, create a plausible question about the video, and provide a short answer with less than three words.
Always answer as if you are directly looking at the video.
[…]
{mdframed}
[align=center,linewidth=1pt]
Question: What did the baby girl on the left side wearing a grey t-shirt do with the egg after she is carrying it?
Answer: The girl throws the egg at the head of the man.
Captions: A boy wearing black shorts is standing on the side of the swimming pool over small rocks and then he performs a backflip and injured himself. In the background, there is a swimming pool, rocks, trees, and people’s voices and water splashing sound is audible.
Question: What was the boy wearing black shorts doing before he performing a backflip?
Answer: The boy is standing on the side of the swimming pool over small rocks. Then he performs a backflip and injured himself.
Captions: A boy wearing a black t-shirt rides a black bicycle in a backward direction and falls on a gray surface. A girl wearing a black cloth is moving on a gray surface, stops, and looks back toward the boy. In the background, there are gray surfaces, buildings, metallic barrier poles, trees, boats, and the sky is visible, and sounds of people speaking and the wind are audible.
Question: What did the girl do after the boy wearing a black t-shirt rides a black bicycle in a backward direction and falls on a gray surface?
Answer: The girl wearing a black cloth is moving on a gray surface. After the boy falls on a gray surface, she stops and looks back toward the boy.
{mdframed}
[align=center,linewidth=1pt]
Captions: A baby girl on the left side wearing a grey t-shirt is carrying an egg then she throws the egg at the head of the man, then the egg falls on the ground and it breaks on a grey surface. A man wearing a red t-shirt sitting on his knees is talking with the baby girl on a grey surface. In the background, there is a grey car, a grey surface, a brown mat, and people speaking and crying sounds are audible.
Question: Why did the egg fall on the ground and breaks?
Answer: The egg was thrown by the girl at the head of the man sitting on his knees. However, it did not hit the man. Therefore, the egg falls on the ground and breaks on a grey surface.
Captions: A boy wearing black shorts is standing on the side of the swimming pool over small rocks and then he performs a backflip and injured himself. In the background, there is a swimming pool, rocks, trees, and people’s voices and water splashing sound is audible.
Question: How did the boy standing on the side of the swimming pool over small rocks injure himself?
Answer: The boy standing on the side of the swimming pool over small rocks perform a backflip. Instead of falling into the water, he hit on the small rocks. That is why he injured himself.
Captions: A boy wearing a black t-shirt rides a black bicycle in a backward direction and falls on a gray surface. A girl wearing a black cloth is moving on a gray surface, stops, and looks back toward the boy. In the background, there are gray surfaces, buildings, metallic barrier poles, trees, boats, and the sky is visible, and sounds of people speaking and the wind are audible.
Question: How did the boy wearing a black t-shirt fall on a gray surface?
Answer: The boy wearing a black t-shirt rides a black bicycle in a backward direction. Riding a black bicycle backward is slow and hard to balance. Also, the boy is not good at riding a black bicycle backward. As a result, he fell on a gray surface.
{mdframed}
[align=center,linewidth=1pt]
Captions: A baby girl on the left side wearing a grey t-shirt is carrying an egg then she throws the egg at the head of the man, then the egg falls on the ground and it breaks on a grey surface. A man wearing a red t-shirt sitting on his knees is talking with the baby girl on a grey surface. In the background, there is a grey car, a grey surface, a brown mat, and people speaking and crying sounds are audible.
Question: who throws the egg at the man
Answer: baby girl
Captions: A boy wearing black shorts is standing on the side of the swimming pool over small rocks and then he performs a backflip and injured himself. In the background, there is a swimming pool, rocks, trees, and people’s voices and water splashing sound is audible.
Question: what kind of pool is in the background
Answer: swimming
Captions: A boy wearing a black t-shirt rides a black bicycle in a backward direction and falls on a gray surface. A girl wearing a black cloth is moving on a gray surface, stops, and looks back toward the boy. In the background, there are gray surfaces, buildings, metallic barrier poles, trees, boats, and the sky is visible, and sounds of people speaking and the wind are audible.
Question: what happens when the man loses control
Answer: falls down